Whats the equation for the line through (-2,4) parallel to the line through )1,1) and (5,7)
Answer by nerdybill(7384)

Whats the equation for the line through (-2,4) parallel to the line through )1,1) and (5,7)
.
Begin by finding slope of:
(1,1) and (5,7)
slope = (y2-y1)/(x2-x1)
slope = (7-1)/(5-1)
slope = 6/4
slope = 3/2
Since the new line is parallel, it must have the same slope of 3/2.
plug that and the given point (-2,4) into the "point-slope" form:
y - y1 = m(x - x1)
y - 4 = (3/2)(x - (-2))
y - 4 = (3/2)(x+2)
y - 4 = (3/2)x + 3
y = (3/2)x + 3 + 4
y = (3/2)x + 7 (this is what they're looking for, in "slope-intercept" form)